Lopez spent 1/3 of his vacation money for travel and 2/5 of his vacation money for lodging. He spent $1100 for travel and lodging. What is the total amount of money he spent on his vacation?
Answer by macston(5194)

.
x=total vacation money
.
(1/3)x+(2/5)x=$1100
(5/15)x+(6/15)x=$1100
(11/15)x=$1100
x=$1500
ANSWER: Lopez spent a total of $1500 on his vacation.